State whether the null hypothesis should be rejected on the basis of given P-value. P-value 0.002, a 0.01, one-tailed test.

1
Expert's answer
2021-12-15T17:10:19-0500

The decision rule is, reject the null hypothesis if "p-value\lt \alpha".

For this case, "p-value=0.002" and "\alpha=0.01".

Since "p-value=0.002\lt \alpha=0.01," we reject the null hypothesis.

Therefore, the decision is to reject the null hypothesis.